# Fine-Grained Cross-View Localization via Local Feature Matching and Monocular Depth Priors

### v1

@@ Appendix E Additional Details on InfoNCE Losses
Appendix F Runtime and Memory Usage
Next, we report the runtime and memory usage of our method on VIGOR.
Runtime: On a single H100 GPU, our method achieves 14.74 FPS without RANSAC Fischler and Bolles (1981) (0.03s for our model and 0.04s for the depth predictor Unik3D Piccinelli et al. (2025)) and 0.42 FPS with RANSAC. This is faster than the previous local feature matching method FG2 Xia and Alahi (2025), which runs at 9.26 FPS without RANSAC and 0.32 FPS with it. Notably, when prioritizing speed, omitting RANSAC results in only a small increase in localization error, e.g., on VIGOR same-area test set with known orientation the mean error increases from 3.06 m to 3.29 m. Furthermore, one can use a faster depth predictor to further reduce inference time.
Memory: our model uses 726.74 MB, with 580.54 MB for the frozen DINOv2 Oquab et al. (2024), while the metric depth predictor Unik3D Piccinelli et al. (2025) uses 1368.81 MB.
Appendix G Limitations
As stated in Sec. 3.2 of the main paper, our method is based on the assumption that the transformation between aerial and ground points can be modeled as a similarity transform. However, aerial images in existing datasets are often not orthorectified. For instance, building facades are visible in some aerial views, indicating perspective distortions. Consequently, the relationship between aerial points and BEV ground points may deviate from a true similarity transformation. Additionally, our pose estimation depends on an external depth predictor. Inaccuracies in the predicted depth maps can directly affect the precision of the estimated poses, potentially leading to degraded localization performance. Still, we experimentally show that, under the assumption of a similarity transformation, using off-the-shelf depth predictors can yield satisfactory localization performance.
Appendix H Broader Impact
Fine-grained cross-view localization has the potential to significantly advance a wide range of real-world applications. In domains such as autonomous driving, robotics, and augmented reality, accurate and reliable localization is essential. Our method offers enhanced robustness and precision in scenarios where specialized sensors are unreliable, such as GNSS in urban canyons. However, we also acknowledge potential risks. Fine-grained localization could be misused for surveillance or unauthorized tracking if deployed without appropriate safeguards. We recommend that future deployments include rigorous ethical guidelines and privacy-preserving mechanisms.

### v2

@@ Appendix G Additional Details on InfoNCE Losses
Appendix H Runtime and Memory Usage
Next, we report the runtime and memory usage of our method on VIGOR in Tab. 11.
Method Runtime (FPS) Memory Usage
No RANSAC With RANSAC
FG2 9.26 0.32 811.52 MB
Ours 14.74 0.42
726.74 MB (base) + 1368.81 MB (Unik3D)
726.74 MB (base) + 117.44 MB (UniFuse)
726.74 MB (base) + 200.23 MB (BiFuse++)
Table 11: Runtime and memory usage comparison.
Runtime: On a single H100 GPU, our method achieves 14.74 FPS without RANSAC (Fischler & Bolles, 1981) (0.03s for our model and 0.04s for the depth predictor Unik3D (Piccinelli et al., 2025)) and 0.42 FPS with RANSAC. This is faster than the previous local feature matching method FG2 (Xia & Alahi, 2025), which runs at 9.26 FPS without RANSAC and 0.32 FPS with it. Notably, when prioritizing speed, omitting RANSAC results in only a small increase in localization error, e.g., on VIGOR same-area test set with known orientation the mean error increases from 3.06 m to 3.29 m. Furthermore, one can use a faster depth predictor to further reduce inference time.
Memory: our model uses 726.74 MB, with 580.54 MB for the frozen DINOv2 (Oquab et al., 2024), while the metric depth predictor Unik3D (Piccinelli et al., 2025) uses 1368.81 MB. Notably, as shown in Tab. 2, our method is also compatible with more lightweight depth models, such as UniFuse Jiang et al. (2021) and BiFuse++ Wang et al. (2022).